OpenStudy (anonymous):

Would ((1+sin(theta))/cos(theta))((1-sin(theta))/cos(theta)) equal 1?

OpenStudy (anonymous):

\[(1+\sin \theta/\cos \theta)(1-\sin \theta/\cos \theta)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I thought so, I just wanted to double check :) Thank you!

OpenStudy (anonymous):

yw:)

OpenStudy (anonymous):

just to confirm your Question was? \[\Large \frac{(1+\sin(\theta))}{\cos(\theta)}*\frac{(1-\sin(\theta))}{\cos(\theta)}\]

OpenStudy (anonymous):

if it was then yes it is equal to 1.